### v1

# Some lemmas on spectral radius of graphs: including an application

Abstract
For a graph $G$, the spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. In this paper, we give three lammas on $\rho(G)$ when $G$ contains a spanning complete bipartite graph. Moreover, an application was also included at the end.
Keywords: spectral radius; walk; cycle; planar graph.
2020 Mathematics Subject Classification: 05C50.
1 Introduction
All graphs considered in this paper are finite, undirected and simple. For a graph $G$, let $\overline{G}$ be its complement. The vertex set and edge set of $G$ are denoted by $V(G)$ and $E(G)$, respectively. For two vertices $u,v$, let $d_{G}(u,v)$ denote the distance between them (i.e., the minimum length of a path connecting them). We say $u\sim v$, if $u$ and $v$ are adjacent in $G$. For a certain integer $n$, let $K_{n},P_{n}$ and $C_{n}$ be the complete graph, the path and the cycle of order $n$, respectively. For any $\ell\geq 2$ graphs $G_{1},G_{2},...,G_{\ell}$, let $\cup_{1\leq i\leq\ell}G_{i}$ be the disjoint union of them. In particular, let $tG=\cup_{1\leq i\leq t}G$ for any integer $t\geq 1$. Let $G_{1}\vee G_{2}$ be the join of $G_{1}$ and $G_{2}$, obtained from $G_{1}\cup G_{2}$ by connecting all vertices of $G_{1}$ to all vertices of $G_{2}$. For any terminology used but not defined here, one may refer to [2].
Let $G$ be a graph with vertices $v_{1},v_{2},...,v_{n}$. The adjacency matrix of $G$ is an $n\times n$ matrix $(a_{ij})$, where $a_{ij}=1$ if $v_{i}\sim v_{j}$, and $a_{ij}=0$ otherwise. The spectral radius $\rho(G)$ of $G$ is the largest eigenvalue of its adjacency matrix. By the well-known Perron–Frobenius theorem, $\rho(G)$ has a non-negative eigenvector. A non-negative eigenvector corresponding to $\rho(G)$ is called a Perron vector of $G$. If $G$ is connected, any Perron vector of $G$ has positive entries. The spectral radius of a graph is closely related to its convergence property. In this paper, we give the following three lammas on spectral radius of graphs, which will be used.
Lemma 1.1
For integers $n_{1}\geq n_{2}+2\geq 1$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|\geq 0$. Then $\rho(G_{1})>\rho(G_{2})$.
Lemma 1.2
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq 1$ and $n_{1}\geq n_{2}+2$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 4}n_{s}\geq 100$. Then $\rho(G_{1})<\rho(G_{2})$.
Lemma 1.3
For integers $n_{1}\geq n_{2}\geq n_{3}\geq n_{4}\geq n_{5}\geq 1$ and $n_{1}\geq n_{2}+3$, let $G_{1}=H\vee(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T)$ and $G_{2}=H\vee(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T)$, where $H,T$ are two graphs with $|H|\geq 1$ and $|T|+\sum_{1\leq s\leq 5}n_{s}\geq 220$. Then $\rho(G_{1})<\rho(G_{2})$.
The rest of the paper is organized as follows. In Section 2, we include several lemmas, which will be used later. The proofs of Lemmas 1.1, 1.2 and 1.3 are given in Section 3. In Section, we include an application on spectral radius of $C_{\ell}$-free planar graphs of order $n\geq 1.8\times 10^{17}$, where $5\leq\ell\leq n$.
2 Preliminaries
Let $G$ be a graph. For an integer $\ell\geq 1$, a walk of length $\ell$ in $G$ is an ordered sequence of vertices $v_{0}v_{1}\cdots v_{\ell}$, such that $v_{i}\sim v_{i+1}$ for any $0\leq i\leq\ell-1$. The vertex $v_{0}$ is called the starting vertex of the walk. For any $u\in V(G)$ and $\ell\geq 1$, let $w^{\ell}_{G}(u)$ be the number of walks of length $\ell$ starting at $u$ in $G$. Let $W^{\ell}(G)=\sum_{v\in V(G)}w^{\ell}_{G}(u)$. The following spectral formula was given in [9].
Theorem 2.1
([9]) For an integer $r\geq 2$, let $K_{n_{1},n_{2},...,n_{r}}$ be the complete $r$-partite graph of order $n$ with parts $V_{1},V_{2},...,V_{r}$, where $n_{i}=|V_{i}|\geq 1$ for $1\leq i\leq r$ and $\sum_{1\leq i\leq r}n_{i}=n$. For each $1\leq i\leq r$, let $H_{i}$ be a graph with vertex set $V(H_{i})\subseteq V_{i}$. Let $G$ be the graph obtained from $K_{n_{1},n_{2},...,n_{r}}$ by embedding the edges of $H_{i}$ into $V_{i}$ for all $1\leq i\leq r$. Set $\rho=\rho(G)$. Then
$\sum_{1\leq s\leq r}\frac{1}{1+\frac{n_{s}}{\rho}+\sum^{\infty}_{i=1}\frac{W^{i}(H_{s})}{\rho^{i+1}}}=r-1.$
The existence of the infinite series in Theorem 2.1 was guaranteed by the following two results. The first one is taken from Theorem 8.1.1 of [2].
Lemma 2.2
([2]) If $H$ is a subgraph of a connected graph $G$, then $\rho(H)\leq\rho(G)$, with equality if and only if $H=G$.
Lemma 2.3
([9]) Let $G$ be a graph of order $n$. Then $\sum^{\infty}_{k=1}\frac{W^{k}(G)}{x^{k}}$ exists for any $x>\rho(G)$.
3 Proofs of Lemmas 1.1, 1.2 and 1.3
Let $G$ be a graph. For a vertex $u$ and a walk $Q$ of $G$, we say that $Q$ crosses $u$, if $u$ is one of the vertices in the sequence of $Q$. For any integer $\ell\geq 1$ and $u,v\in V(G)$, let $W^{\ell}_{u,v}(G)$ be the number of walks $Q$ of length $\ell$ in $G$, such that $Q$ crosses $u$ and $v$. Let $P_{0}$ denote the empty graph.
Lemma 3.1
For integers $n_{1}\geq n_{2}+2\geq 3$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. Then $W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$ for any $\ell\geq 1$.
Proof: Let $G_{1}=P_{n_{1}}\cup P_{n_{2}}$ and $G_{2}=P_{n_{1}-1}\cup P_{n_{2}+1}$. For any given $\ell\geq 1$, we shall prove
$W^{\ell}(G_{1})-W^{\ell}(G_{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}).$
Recall that $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$ in $G_{1}$. In $G_{2}$, denote $P_{n_{2}+1}=v_{1}v_{2}\cdots v_{n_{2}}v_{n_{2}+1}$. Let $X$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$. Let $Y$ denote the set of walks $Q$ of length $\ell$ in $G_{2}$, such that $Q$ crosses $v_{1}$. Clearly, $W^{\ell}(G_{1})=|X|+W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}})$ and $W^{\ell}(G_{2})=|Y|+W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}})$. This means that
$W^{\ell}(G_{1})-W^{\ell}(G_{2})=|X|-|Y|.$
Now we divide $X$ in two parts: let $X_{1}$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$ and $u_{n_{2}+2}$; let $X_{2}$ denote the set of walks $Q$ of length $\ell$ in $G_{1}$, such that $Q$ crosses $u_{1}$ and does not cross $u_{n_{2}+2}$. Clearly, $X=X_{1}\cup X_{2}$. By definitions of the sets $X_{1},X_{2},Y$, we see $|X_{2}|=|Y|$ and $|X_{1}|=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. Thus $|X|-|Y|=|X_{1}|=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. Hence, $W^{\ell}(G_{1})-W^{\ell}(G_{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$. This completes the proof. $\Box$
Lemma 3.2
Let $G$ be a path of order $n\geq 3$. Let $u,v$ be two vertices of $G$ such that $d_{G}(u,v)=\ell$, where $2\leq\ell\leq n-1$. Then $\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4})$ for any $x\geq\max\left\{\sqrt{n},5\right\}$.
Proof: Let $P_{\infty}$ be a path of ”infinite” order, such that $u,v$ are two vertices of distance $\ell$ in $P_{\infty}$. That is,
$P_{\infty}=\cdots u_{\ell}\cdots u_{2}u_{1}uz_{1}z_{2}\cdots z_{\ell-1}vv_{1}v_{2}\cdots.$
(See Figure 1, where the ”semicircle” denotes the distance $\ell$ between $u$ and $v$.) Clearly, $W^{s}_{u,v}(G)\leq W^{s}_{u,v}(P_{\infty})$ for any $s\geq 1$. For any vertex $y\in P_{\infty}$ and integer $a\geq 1$, let $w^{a}_{uv}(y)$ be the number of walks $Q$ of length $a$ in $P_{\infty}$, such that $Q$ starts at $y$ and crosses $u$ and $v$.
Figure 1: The infinite path $P_{\infty}$
Claim 1. $w^{a}_{uv}(z_{i})\leq w^{a}_{uv}(u_{i})+w^{a}_{uv}(v_{\ell-i})$ for any $1\leq i\leq\ell-1$ and $a\geq 1$.
Proof of Claim 1. Let $a\geq 1$ be given. Clearly, $w^{a}_{uv}(z_{i})=0$ for $1\leq a\leq\ell-1$. Now assume $a\geq\ell$. Let $1\leq i_{0}\leq\ell-1$ be fixed. We will show that
$w^{a}_{uv}(z_{i_{0}})\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}}).$
Let $X$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u$ first before $v$. Let $Y$ be the set of walks $Q$ starting at $z_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $v$ first before $u$. Clearly, $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|$.
Now we show that $|X|\leq w^{a}_{uv}(u_{i_{0}})$ and $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. We show the first one. Let $D$ be the set of walks $Q$ starting at $u_{i_{0}}$ of length $a$ in $P_{\infty}$, such that $Q$ crosses $u,v$. We can define an injective map $\phi$ between $X$ and $D$. For any $Q\in X$, since $Q$ crosses $u$ first, we can denote
$Q=z_{i_{0}}z_{j_{1}}\cdots z_{j_{s}}uf_{1}f_{2}\cdots,$
where $1\leq j_{1},...,j_{s}\leq\ell-1$ and $f_{1},f_{2},$... are vertices of $P_{\infty}$. Define
$\phi(Q)=u_{i_{0}}u_{j_{1}}\cdots u_{j_{s}}uf_{1}f_{2}\cdots.$
It is easy to see that $\phi$ is injective. Hence, $|X|\leq|D|=w^{a}_{uv}(u_{i_{0}})$. Similarly, we can show that $|Y|\leq w^{a}_{uv}(v_{\ell-i_{0}})$. Then $w^{a}_{uv}(z_{i_{0}})=|X|+|Y|\leq w^{a}_{uv}(u_{i_{0}})+w^{a}_{uv}(v_{\ell-i_{0}})$. This finishes the proof of Claim 1. $\Box$
By Claim 1, we see
$\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})\leq\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(v_{i})$
for any $a\geq 1$. Clearly, $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})=0$ for any $i\geq\ell+1$ and $a\leq 2\ell$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})$ $=w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell-1}w^{a}_{uv}(z_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})$
$\leq 2(w^{a}_{uv}(u)+w^{a}_{uv}(v)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})+\sum_{1\leq i\leq\ell}w^{a}_{uv}(v_{i})).$
By symmetry, $w^{a}_{uv}(u)=w^{a}_{uv}(v)$ and $w^{a}_{uv}(u_{i})=w^{a}_{uv}(v_{i})$ for any $a\geq 1$ and $i\geq 1$. Thus, for any $a\leq 2\ell$,
$W^{a}_{u,v}(P_{\infty})\leq 4(w^{a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{a}_{uv}(u_{i})).$
Since $P_{\infty}$ is a path, each walk starting at $u_{i}$ corresponds to a $\pm$”-series:
$u_{i}+--+-+\cdots.$
(For example, $u_{1}+-++$-- corresponds to the walk $u_{1}uu_{1}uz_{1}uu_{1}$.) For any $a\geq 0$ and $1\leq i\leq a$, in the corresponding $\pm$”-series to the walks of length $\ell+a$ starting at $u_{i}$ and crossing $u$ and $v$, there are at most $\lfloor\frac{a-i}{2}\rfloor$ symbols ”-” (in the all $\ell+a$ symbols $\pm$”). Thus,
$w^{\ell+a}_{uv}(u_{i})\leq\binom{\ell+a}{\lfloor\frac{a-i}{2}\rfloor}\leq\frac{(\ell+a)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}.$
Hence, $w^{\ell+a}_{uv}(u_{i})\leq\frac{(2\ell)^{\lfloor\frac{a-i}{2}\rfloor}}{\lfloor\frac{a-i}{2}\rfloor!}$ for any $0\leq a\leq\ell$. It follows that, for any $0\leq a\leq\ell$,
$w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i})\leq 2\times(1+2\ell+\frac{(2\ell)^{2}}{2!}+\cdots+\frac{(2\ell)^{\lfloor\frac{a}{2}\rfloor}}{\lfloor\frac{a}{2}\rfloor!})=2\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}.$
Then
$\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}$ $\leq\sum_{0\leq a\leq\ell}\frac{4}{x^{\ell+a}}(w^{\ell+a}_{uv}(u)+\sum_{1\leq i\leq\ell}w^{\ell+a}_{uv}(u_{i}))$
$\leq\frac{8}{x^{\ell}}\sum_{0\leq a\leq\ell}\frac{1}{x^{a}}\sum_{0\leq i\leq\lfloor\frac{a}{2}\rfloor}\frac{(2\ell)^{i}}{i!}$
$=\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\sum_{2i\leq a\leq\ell}\frac{1}{x^{a}}$
$\leq\frac{8}{x^{\ell}}\sum_{0\leq i\leq\lfloor\frac{\ell}{2}\rfloor}\frac{(2\ell)^{i}}{i!}\frac{1}{x^{2i}}\frac{x}{x-1}$
$\leq\frac{8}{x^{\ell}}\frac{x}{x-1}\sum^{\infty}_{i=0}\frac{1}{i!}(\frac{2\ell}{x^{2}})^{i}$
$=\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
Note that $W^{s}_{u,v}(G)=0$ for any $s\leq\ell-1$ as $d_{G}(u,v)=\ell$. Then
$\sum_{1\leq s\leq 2\ell}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\sum_{0\leq a\leq\ell}\frac{W^{\ell+a}_{u,v}(P_{\infty})}{x^{\ell+a}}\leq\frac{8}{x^{\ell}}\frac{x}{x-1}e^{\frac{2\ell}{x^{2}}}.$
Since $G$ has maximum degree 2, we see $W^{a}_{u,v}(G)\leq 2^{a}n$ for any $a\geq 1$. Noting $s-3\leq 2(s-\ell-2)$ for $s\geq 2\ell+1$, we have
$\sum_{s\geq 2\ell+1}\frac{W^{s}_{u,v}(G)}{x^{s}}$ $\leq\sum_{s\geq 2\ell+1}\frac{2^{s}n}{x^{s}}$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}\frac{2^{s-3}}{x^{s-\ell-2}}~~~(using~x\geq\sqrt{n})$
$\leq\frac{8}{x^{\ell}}\sum_{s\geq 2\ell+1}(\frac{4}{x})^{s-\ell-2}~~~(using~s-3\leq 2(s-\ell-2))$
$\leq\frac{8}{x^{\ell}}\sum_{j\geq 1}(\frac{4}{x})^{j}~~~(using~\ell\geq 2)$
$=\frac{1}{x^{\ell}}\frac{32}{x-4}~~~(using~x\geq 5).$
Consequently,
$\sum_{s\geq 1}\frac{W^{s}_{u,v}(G)}{x^{s}}\leq\frac{1}{x^{\ell}}(\frac{8x}{x-1}e^{\frac{2\ell}{x^{2}}}+\frac{32}{x-4}).$
This completes the proof. $\Box$
Now we are ready to prove the three lemmas.
Proof of Lemma 1.1. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup T$ and $G^{2}=P_{n_{1}-1}\cup P_{n_{2}+1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}>\rho_{2}$. By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
(Clearly, the domain of $f(G_{j},x)$ is of the form $x\geq a_{j}$ for some positive number $a_{j}$.) Clearly, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain. Thus $\rho_{j}$ is the largest root of $f(G_{j},x)=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}).$
This means that $W^{\ell}(G^{1})-W^{\ell}(G^{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\geq 0$. Clearly, $W^{n_{2}+1}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})=2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{2}^{i}}>\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{2}^{i}}.$
This implies that
$f(G_{1},\rho_{2})<f(G_{2},\rho_{2}).$
Note that $f(G_{2},\rho_{2})=1$. So, $f(G_{1},\rho_{2})<1$. Recall that $f(G_{1},x)$ is strictly increasing in its domain, and $f(G_{1},\rho_{1})=1$. We must have $\rho_{1}>\rho_{2}$. This completes the proof. $\Box$
Proof of Lemma 1.2. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup T$ and $G^{2}=P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}-1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}<\rho_{2}$. Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}}$ as a subgraph, we have
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 4}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 4}n_{i}}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}.$
By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
Similar to Lemma 1.1, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain, and $f(G_{j},\rho_{j})=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. In $G^{2}$, denote $P_{n_{3}+1}=v_{1}v_{2}\cdots v_{n_{3}+1}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}}),$
and
$W^{\ell}(P_{n_{3}+1}\cup P_{n_{4}-1})-W^{\ell}(P_{n_{3}}\cup P_{n_{4}})=W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
This means that
$W^{\ell}(G^{1})-W^{\ell}(G^{2})=W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
Note that $W^{i}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\leq W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})$ for any $i\geq 1$ by the definition of $W^{i}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})$, since $n_{2}\geq n_{4}$. Thus,
$W^{\ell}(G^{1})-W^{\ell}(G^{2})\leq W^{\ell}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1}).$
Then
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})=\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})-W^{i}(G^{2})}{\rho_{1}^{i}}\leq\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})-W^{i}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})}{\rho_{1}^{i}}.$
Clearly, $W^{n_{4}}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})=2$ and $W^{n_{4}+1}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})\geq 2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}_{v_{1},v_{n_{4}+1}}(P_{n_{3}+1})}{\rho_{1}^{i}}\geq\frac{2}{\rho_{1}^{n_{4}}}+\frac{2}{\rho_{1}^{n_{4}+1}}=\frac{1}{\rho_{1}^{n_{4}+1}}(2\rho_{1}+2).$
By Lemma 3.2 (by letting $\ell=n_{4}+1$ there),
$\sum^{\infty}_{i=1}\frac{W^{i}_{u_{1},u_{n_{4}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{1}{\rho_{1}^{n_{4}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{10,\sqrt{\sum_{1\leq i\leq 4}n_{i}}\right\}\geq\max\left\{10,\sqrt{4n_{4}+2}\right\}$, we have $e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}<e^{0.52}<1.7$. Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
$\leq\frac{2}{\rho_{1}^{n_{4}+1}}(\frac{4\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{4}+1)}{\rho_{1}^{2}}}+\frac{16}{\rho_{1}-4}-\rho_{1}-1)$
[…]
Now consider $g(x)=5.8+\frac{6.8}{x-1}+\frac{16}{x-4}-x$ for $x\geq 10$. Clearly, $g(x)$ is decreasing for $x\geq 10$. By a calculation, $g(10)=-\frac{7}{9}<0$. Thus $g(x)<0$ for any $x\geq 10$. Then $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
Proof of Lemma 1.3. Let $G^{1}=P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}}\cup P_{n_{4}}\cup P_{n_{5}}\cup T$ and $G^{2}=P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1}\cup P_{n_{4}+1}\cup P_{n_{5}-1}\cup T$. Then $G_{j}=H\vee G^{j}$ for $j=1,2$. Set $\rho(G_{j})=\rho_{j}$ for any $1\leq j\leq 2$. We will prove $\rho_{1}<\rho_{2}$. Since $G_{j}$ contains $K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}}$ as a subgraph, we have
$\rho_{j}\geq\rho(K_{1}\vee\overline{K_{|T|+\sum_{1\leq i\leq 5}n_{i}}})=\sqrt{|T|+\sum_{1\leq i\leq 5}n_{i}}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}.$
By Theorem 2.1,
$\frac{1}{1+\frac{|H|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{\rho^{i}_{j}}}+\frac{1}{1+\frac{|G^{j}|}{\rho_{j}}+\frac{1}{\rho_{j}}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{\rho^{i}_{j}}}=1,$
where $1\leq j\leq 2$. For $1\leq j\leq 2$, let
$f(G_{j},x)=\frac{1}{1+\frac{|H|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(H)}{x^{i}}}+\frac{1}{1+\frac{|G^{j}|}{x}+\frac{1}{x}\sum^{\infty}_{i=1}\frac{W^{i}(G^{j})}{x^{i}}}.$
Clearly, $f(G_{j},x)$ is strictly increasing with respective to $x$ in its domain, and $f(G_{j},\rho_{j})=1$.
In $G^{1}$, denote $P_{n_{1}}=u_{1}u_{2}\cdots u_{n_{1}}$. In $G^{2}$, denote $P_{n_{4}+1}=v_{1}v_{2}\cdots v_{n_{4}+1}$. For any $\ell\geq 1$, by Lemma 3.1,
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}})=W^{\ell}(P_{n_{1}}\cup P_{n_{2}})-W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1})$
$W^{\ell}_{u_{1},u_{n_{2}+2}}(P_{n_{1}})\leq W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
Similarly,
$W^{\ell}(P_{n_{1}-1}\cup P_{n_{2}+1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1})=W^{\ell}(P_{n_{1}-1}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{3}+1})$
$W^{\ell}_{u_{1},u_{n_{3}+2}}(P_{n_{1}-1})\leq W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
It follows that
$W^{\ell}(P_{n_{1}}\cup P_{n_{2}}\cup P_{n_{3}})-W^{\ell}(P_{n_{1}-2}\cup P_{n_{2}+1}\cup P_{n_{3}+1})\leq 2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}}).$
Also by Lemma 3.1,
$W^{\ell}(P_{n_{4}+1}\cup P_{n_{5}-1})-W^{\ell}(P_{n_{4}}\cup P_{n_{5}})=W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1}).$
It follows that
$W^{\ell}(G^{1})-W^{\ell}(G^{2})\leq 2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1}).$
Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})\leq\sum^{\infty}_{i=1}\frac{2W^{\ell}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})-W^{\ell}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})}{\rho_{1}^{i}}.$
Clearly, $W^{n_{5}}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})=2$ and $W^{n_{5}+1}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})\geq 2$. It follows that
$\sum^{\infty}_{i=1}\frac{W^{i}_{v_{1},v_{n_{5}+1}}(P_{n_{4}+1})}{\rho_{1}^{i}}\geq\frac{2}{\rho_{1}^{n_{5}}}+\frac{2}{\rho_{1}^{n_{5}+1}}=\frac{1}{\rho_{1}^{n_{5}+1}}(2\rho_{1}+2).$
By Lemma 3.2 (by letting $\ell=n_{5}+1$ there),
$\sum^{\infty}_{i=1}\frac{2W^{i}_{u_{1},u_{n_{5}+2}}(P_{n_{1}})}{\rho_{1}^{i}}\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}).$
Since $\rho_{1}\geq\max\left\{\sqrt{220},\sqrt{\sum_{1\leq i\leq 5}n_{i}}\right\}\geq\max\left\{\sqrt{220},\sqrt{5n_{5}+3}\right\}$, we have $e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}<e^{0.405}<1.5$. Thus,
$\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})$
$\leq\frac{2}{\rho_{1}^{n_{5}+1}}(\frac{8\rho_{1}}{\rho_{1}-1}e^{\frac{2(n_{5}+1)}{\rho_{1}^{2}}}+\frac{32}{\rho_{1}-4}-\rho_{1}-1)$
[…]
Now consider $g(x)=11+\frac{12}{x-1}+\frac{32}{x-4}-x$ for $x\geq\sqrt{220}$. Clearly, $g(x)$ is decreasing for $x\geq\sqrt{220}$. By a calculation, $g(\sqrt{220})=-0.01\cdots$. Thus $g(x)<0$ for any $x\geq\sqrt{220}$. So, $\sum^{\infty}_{i=1}(\frac{W^{i}(G^{1})}{\rho_{1}^{i}}-\frac{W^{i}(G^{2})}{\rho_{1}^{i}})<0$, or equivalently, $\sum^{\infty}_{i=1}\frac{W^{i}(G^{1})}{\rho_{1}^{i}}<\sum^{\infty}_{i=1}\frac{W^{i}(G^{2})}{\rho_{1}^{i}}$. It follows that $f(G_{1},\rho_{1})>f(G_{2},\rho_{1})$. Note that $f(G_{1},\rho_{1})=1$. So, $f(G_{2},\rho_{1})<1$. Recall that $f(G_{2},x)$ is strictly increasing in its domain, and $f(G_{2},\rho_{2})=1$. We must have $\rho_{1}<\rho_{2}$. This completes the proof. $\Box$
4 Spectral radius of $C_{\ell}$-free planar graphs
A graph is called planar, if it can be drawn in the plane without crossed edges. Recently, Tait and Tobin [6] proved that $K_{2}\vee P_{n-2}$ is the only extremal graph with the maximum spectral radius among all planar graphs of large order. This (asymptotically) solves a conjecture proposed by Boots and Royle [1], and independently by Cao and Vince [3]. It is also interesting for the researchers to determine the extremal graphs with the maximum spectral radius among all $C_{\ell}$-free planar graphs of order $n$, where $3\leq\ell\leq n$. The case $\ell=4$ was settled by Nikiforov [5], and Zhai and Wang [8]. Very recently, Fang, Lin and Shi [4] solved the case when $\ell=3$ and $\ell\ll n$. Xu, Lin and Fang [7] solved the case when $\frac{\ell}{n}$ is close to 1. They [7] proposed it as a problem for the general case. Armed with Lemma 1.1 and Lemma 1.2, using the typical spectral method we can solve this problem. That is, we have the following Theorem 4.1.
For a finite family of graphs $\mathcal{G}$, let ${\rm SPEX}(\mathcal{G})$ be the set of extremal graphs with the maximum spectral radius in $\mathcal{G}$.
Theorem 4.1
For integers $n\geq 1.8\times 10^{17}$ and $5\leq\ell\leq n$, let $\mathcal{G}_{n,\ell}$ be the set of all $C_{\ell}$-free planar graphs of order $n$. Then we have the following conclusions.
$(i)$ For $5\leq\ell<\frac{2n+5}{3}$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{\lceil\frac{\ell-3}{2}\rceil}\cup(a+1)P_{\lfloor\frac{\ell-3}{2}\rfloor}\cup P_{b})\right\}$, where $n-\ell+1=a\lfloor\frac{\ell-3}{2}\rfloor+b$ with $0\leq b<\lfloor\frac{\ell-3}{2}\rfloor$.
$(ii)$ For $\frac{2n+5}{3}\leq\ell\leq n$, ${\rm SPEX}(\mathcal{G}_{n,\ell})=\left\{K_{2}\vee(P_{2\ell-n-4}\cup 2P_{n-\ell+1})\right\}$.
Declaration of competing interest
There is no conflict of interest.
Data availability statement
No data was used for the research described in the article.